OpenStudy (anonymous):

Proving definite integral Trying to prove that \[\frac{ 1 }{ \sqrt \pi } \int\limits_{-\infty}^{\infty} e^{-y^2} dy = 1\] This actually stems from an electromagnetism question solving the delta function, however I've gotten this far and now am drawing a blank as to how to prove this? I 've used substitution to obtain this simplified form and am not sure where to go from here?

OpenStudy (anonymous):

I have now reached the same point with a different delta funcion, obtaining \[ \int\limits_{- \infty}^{\infty} \frac{ 1 }{ \pi } \frac{ \sin y }{ y} dy = 1\] I have tried to use integration by parts for this but I seem to be stuck at the evaluation part? Any help would be greatly appreciated

OpenStudy (anonymous):

substitute y^2=t and use the properties of even function prior to it...

OpenStudy (ankitshaw):

siny/y is not integratable

zepdrix (zepdrix):

I remember showing something similar to this using the Gamma Function. Lemme keep trying, I think I've almost got it :U

OpenStudy (anonymous):

@ankitshaw It definitely does work, http://www.wolframalpha.com/input/?i=%5Cint%5Climits_%7B-+%5Cinfty%7D%5E%7B%5Cinfty%7D+%5Cfrac%7B+1+%7D%7B+%5Cpi+%7D+%5Cfrac%7B+%5Csin+y+%7D%7B+y%7D+dy+ I'm just not sure now, hanks @zepdrix :-)

zepdrix (zepdrix):

Grr I feel like I'm getting close :[ But there are still some holes in my steps... Lemme show you what I was trying,

zepdrix (zepdrix):

Starting with this, \[\Large\rm \int\limits\limits_{-\infty}^{\infty} e^{-y^2} dy\]Making the substitution,\[\Large\rm u=y^2\]\[\Large\rm du=2y~dy\]And then you have to do some fancy jazz to get it in terms of u, Oh let's uhhhh split the original integral in half, that might solve the problem I was running into,\[\Large\rm \int\limits_{-\infty}^{0} e^{-y^2}dy+\int\limits_{0}^{\infty} e^{-y^2}dy\]And then use rules of integration,\[\Large\rm -\int\limits\limits_{0}^{-\infty} e^{-y^2}dy+\int\limits\limits_{0}^{\infty} e^{-y^2}dy\]So after some fancy foot work you get it in terms of u,\[\Large\rm -\frac{1}{2}\int\limits_0^{\infty}e^{-u}u^{-1/2}~du+\frac{1}{2}\int\limits_0^{\infty}e^{-u}u^{-1/2}~du\]Hmm but then they end up canceling out :-( Yah I made a boo boo somewhere in there... Anyway, I'm trying to write it in this form:\[\Large\rm \Gamma(t)=\int\limits_0^{\infty}e^{-u}u^{t-1}~du\]So we would just change the -1/2 to 1/2-1 which let's us know we're dealing with Gamma of 1/2.\[\Large\rm \Gamma(1/2)\]Which is equal to sqrt(pi). (We would have to go through a bit of a longer process if we needed to prove that though).

zepdrix (zepdrix):

bah i dunno :c i gotta get some sleep right now

OpenStudy (anonymous):

Try to see an easy proof using polar coordinates and double integration; The details are posted in http://en.wikipedia.org/wiki/Gaussian_integral

OpenStudy (primeralph):

Yeah the methods described are right, but the polar one is what's used often. It's easier.

OpenStudy (primeralph):

You can also use the Laplace Transform (or Fourier Transform if you have the time).

OpenStudy (primeralph):